The takeoff speed for this plane will be 70 m/s. Assuming this minimum acceleration, what is the minimum

allowed length for the runway?


1
Expert's answer
2020-11-25T10:45:20-0500

Taking the minimum acceleration to be 5m/s2


Given the initial velocity, u=0m/s and final velocity to be v=70m/s.


Acceleration =5m/s2



If "d" be the allowed length of the runway.


Using the formula ;"v^{2} - u^{2} =2aS"


"=(70) ^{2} - (0)^{2} =2(5)S"


4900=10S


"S=\frac{4900}{10}"


"d=490m"